Bobby has a pile of dimes and quarters from his coin
jar.
a. In total, Bobby has $24.90. Write an equation in
which the number of quarters, q, and number of
dimes, d, add up to $24.90.
b. If Bobby had 15 more dimes he would the same
amount of dimes as triple the amount of his
quarters. Write another equation for this
situation with q representing the number of
quarters that Bobby has and d representing the
number of dimes he has.
c. Use substitution to solve this system of
equations for the number of quarters and dimes
Bobby has.
d. Explain how you would check your answer in c.
e. Graph your equations in part a) and b). Be sure
you show the point of intersection. What does
this point of intersection mean in the context of
this problem?

Answer by stanbon(75887)

Bobby has a pile of dimes and quarters from his coin
jar.
a. In total, Bobby has $24.90. Write an equation in
which the number of quarters, q, and number of
dimes, d, add up to $24.90.
value + value = 2490 cents
25q + 10d = 2490
------------------------------------------------------------
b. If Bobby had 15 more dimes he would the same
amount of dimes as triple the amount of his
quarters. Write another equation for this
situation with q representing the number of
quarters that Bobby has and d representing the
number of dimes he has.
-------
d+15 = 3q
------------------------------------------------------------
c. Use substitution to solve this system of
equations for the number of quarters and dimes
Bobby has.
25q + 10(3q-15) = 2490
25q + 30q - 150 = 2490
55q = 2640
q = 48
----
d = 3q-15 = 3*48-15 = 129
---------------
Cheers,
Stan H.
-------------
d. Explain how you would check your answer in c.
e. Graph your equations in part a) and b). Be sure
you show the point of intersection. What does
this point of intersection mean in the context of
this problem?